In 9 ​years, a child will be 44 years older than twice her current age. What is the current age of the​ child?
Answer by ankor@dixie-net.com(22740)

In 9 ​years, a child will be 4 years older than twice her current age.
What is the current age of the​ child? (edited to make sense)
:
let x = her current age
x + 9 = 2x + 4
x - 2x = 4 - 9
-x = -5
x = 5 yrs is her current age